OpenStudy (anonymous):

how do u solve 25y^2 - 49w^2

OpenStudy (lgbasallote):

are you familiar with the difference of squares? \[\huge a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

a little yes

OpenStudy (lgbasallote):

good. so what is \(\sqrt{25y^2}\)

OpenStudy (anonymous):

5y... idk

OpenStudy (lgbasallote):

yes it's 5y

OpenStudy (lgbasallote):

now what is \(\sqrt{49w^2}\)

OpenStudy (anonymous):

7w

OpenStudy (lgbasallote):

right. so... a = 5y b = 7w substitute that into (a+b)(a-b) what do you get?

OpenStudy (anonymous):

do you do foil?

OpenStudy (lgbasallote):

no.. just put 5y into a and 7w into b in (a+b)(a-b)

OpenStudy (anonymous):

and thats your whole answer?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

oh... i feel stupid! :) thanks!

OpenStudy (lgbasallote):

welcome